A race car starts from rest on a circular track of radius 400 m. Its speed increases at the constant rate of 0.500 m/s2. At the point where the magnitudes of the radial acceleration is twice the tangential acceleration, determine (a) the speed of the race car, and (b) the elapsed time.


1
Expert's answer
2021-05-03T07:09:39-0400

a)


"2a_t=\frac{v^2}{r}\\2(0.5)=\frac{v^2}{400}\\v=20\frac{m}{s}"

b)


"t=\frac{v}{a}=\frac{20}{0.5}=40\ s"